OpenStudy (anonymous):

Find the product of (x + 5)2. x2 − 10x + 25 x2 − 25 x2 + 25 x2 + 10x + 25

OpenStudy (mayankdevnani):

\[\huge \bf (x+5)^2\] use \[\huge \bf (a+b)^2~identity\] and this identity states that:- \[\huge \bf (a+b)^2=a^2+b^2+2ab\] \[\huge \bf where a=x ~and ~b=5\]

OpenStudy (mayankdevnani):

can you solve it? @supanick14

OpenStudy (anonymous):

You aren't solving it you are multiplying it or simplifying it. It sounds more like you want someone to answer the question for you rather than learn to do it yourself. Are you on here to just "get answers" or learn how to do it?

OpenStudy (skullpatrol):

Failure to follow the guidelines may lead to your suspension or banning from the site. If you see any violations, please report them by using a report abuse link. OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.

OpenStudy (anonymous):

@skullpatrol thanks, will do!